OpenStudy (anonymous):

how to find two numbers which differ by 4 and whose squares differ by 72

OpenStudy (anonymous):

1. b - a = 4 (two numbers differ by 4) 2. b^2 - a^2 = 72 (their squares differ by 72) solve the first one for b: 3. b = a + 4 substitute "a + 4" for "b" in the second one: (a + 4)^2 - a^2 = 72 a^2 + 8a + 16 - a^2 = 72 8a + 16 = 72 8a = 56 a = 7 substitute "7" for "a" in equation 3: b = 7 + 4 b = 11

OpenStudy (anonymous):

yay! thank you!! :)